OpenStudy (anonymous):

will some one please help me im really bad in math. if you can answer my question would you also explain how you got the answer. I will also give you a medal. write y(-3/4)x+3 in standard form using integers (same for this one) y=-1/2x+1

OpenStudy (anonymous):

Standard form of equation is Ax +By = C

OpenStudy (anonymous):

so what does a,b,and c stand for?

OpenStudy (anonymous):

Number (in this case integers) coefficient of x, y and C =constant

OpenStudy (anonymous):

ok but im still not understanding how to do the equation

OpenStudy (anonymous):

Put your equations in Ax+By = c form A, B and C should be integers

OpenStudy (anonymous):

kinda like ok -3x+4y= and then what ever c is??? I really don't get this and my answer choices don't make it any easier

OpenStudy (anonymous):

-3x +4y =-12

OpenStudy (anonymous):

ok how did you get -12?

OpenStudy (anonymous):

3(-4)

OpenStudy (anonymous):

y -(3/4)x +3 = 0 multiply both side by 4 4y -3x +12 =0 -3x +4y = -12

OpenStudy (anonymous):

ok thank you so much so on the second one it would be kinda the same but like y=-1/2x+1 multiply 2 on both sides or is it different?

OpenStudy (anonymous):

Its same.

OpenStudy (anonymous):

ok just looked at my answer choices again and one the first question there isn't a -3x+4y=-12 but there is a 3x+4y=-12 and a -3+4y=12 im not sure about which one to put as my answer and the second one I did what you said and I got x-2y=2 will you see if im right please

OpenStudy (anonymous):

I meant to say the first answer choice doesn't have -3x+4y=-12

OpenStudy (anonymous):

you are correct. its just x+2y = 2 and first one can also be written as 3x -4y = 12

OpenStudy (anonymous):

Thank you so much and yay I got one right lol thank you again

OpenStudy (anonymous):

Anytime :)